# Theories of synaptic memory consolidation and intelligent plasticity for continual learning

### v1

@@ 1 Introduction
The vast neuronal networks that constitute our brain are the substrate underlying this remarkable ability of memory formation and storage. Memories are stored as synaptic efficacy changes, i.e., the strength of the synaptic connections between neurons. Jointly, these changes cause lasting changes to the computational transformation that a given neuronal circuit performs, and it is this change that lies at the heart of memory. For instance, consider a neuronal circuit whose task is to associate a particular reward, e.g., some food, with a specific sensory stimulus. Upon experiencing a certain stimulus, e.g., a tone of a given frequency, the circuit in question may not produce any output initially. That is to say that it has not yet associated any value with this stimulus. Only once the tone is paired consistently with a reward does the brain circuit start producing an output specific to the paired stimulus. Thus, the next time the circuit receives the tone stimulus, its output will signal the value, e.g., the expected reward, even before the food has arrived. Meanwhile, the circuit response to another control stimulus, i.e., another tone, is unchanged. The circuit described above stores a memory of the rewarded stimulus. The paradigm described above is called Pavlovian conditioning, which honors Pavlov, who first performed such an experiment on his dog. And importantly, there are several brain areas, like, for instance, the amygdala, that are involved in such value learning.
@@
A plethora of bio-chemical mechanisms exist that can change synaptic connections in an activity-dependent manner. These mechanisms are jointly called synaptic plasticity and are presumed to be the bio-physical basis of memory and learning. This book covers synaptic plasticity in-depth, so we won’t delve into them too deeply. For now, let us focus on the conceptual challenges these synaptic plasticity mechanisms must overcome to implement memory. There are several challenges. First, the synaptic changes must be stimulus-specific to form a distinct memory. Second, the changes need to be orchestrated so that the circuit does not succumb to pathological activity but remains in a healthy activity regime. Thus, plasticity also needs to ensure the stability of the circuit. This function is typically associated with various forms of homeostatic plasticity in neuronal circuits that act on different length- and timescales. Finally, plasticity must ensure that the circuit remains plastic for storing future memories without compromising the information already stored in the circuit.
@@
Our ability to learn continually is pervasive and intrinsically linked to how we learn. It applies to most learning problems requiring memory. Language is only one example. Other examples are physical skills such as riding a bicycle or playing volleyball. We must learn both skills in an often painstaking process, and they require overlapping sets of muscles, so at least some of the neuronal circuits involved are bound to be shared among them. We learn both tasks at different times, and yet learning to play volleyball does not impair our ability to ride a bike. This chapter will give us an overview of theoretical work on continual learning and memory in neural network models.
We first review the classic Hopfield network, an influential associative memory model, which allows us to define essential concepts such as memory capacity and interference. The latter is linked to forgetting, and we will see how. Equipped with these concepts, we will explore several strands of research suggesting that synapses must possess an internal state or memory in their own right and intelligently use it to store new memories. To understand why, we will get to know the notion of memory lifetime and the fundamental requirement for memory consolidation mechanisms in the continual learning setting. Finally, we will turn to the more contemporary deep artificial neural network (ANN) models, the problem of catastrophic forgetting, and putative solutions based on synaptic consolidation and “intelligent” synaptic plasticity rules.
@@ 2 Learning and forgetting in Hopfield networks
which stores $p$ randomly and independently drawn binary patterns $\xi\in\{-1,1\}^{n}$ in which a one appears with a probability of 0.5. If we look at the above expression, we see that each term of the sum adds positive connections between neurons that are active (+1) in the same memory together. Hence, we can think of learning in the Hopfield model as a Hebbian learning rule which defines an assembly of strongly connected neurons as an engram for each stored memory (Hebb, 1949; cf. Fig. 1).
@@ 2.1 Memory capacity and forgetting
In the Hopfield model, the phenomenon is called “catastrophic forgetting.” Here, a word of caution is in order, as it is the first time we encounter the term catastrophic forgetting, and we will re-encounter it when discussing learning in ANNs (Section 4). We highlight it here because the term refers to fundamentally different phenomena in these two contexts. In the Hopfield model, catastrophic forgetting is linked to its finite memory capacity. It describes a phenomenon that impairs the network’s ability to recall any stored memory patterns through its dynamics. In the context of ANNs, as we will see later, it refers to old memories being quickly overwritten by new memories, a feature also called the palimpsest property. What adds to the confusion is that overwriting old memories through new memories is the solution to avoiding catastrophic forgetting in the Hopfield model. At the same time, forgetting is the primary problem for ANNs. While the whole meaning of this statement will only become clear at the end of this chapter, let us first understand why overwriting old memories solves the issue of catastrophic forgetting in the Hopfield network.
@@ 2.2 The stability-plasticity dilemma
It is called a paradox because the above statement contradicts our daily experience by which we constantly and quickly form new memories. For instance, we readily remember a person to whom we are introduced. But we also retain old memories, some of them for life. Our brains have evolved to deal with this plasticity-stability dilemma very efficiently. They learn continually by adding new memories with minimal impact on many old memories. The underlying synaptic and neuronal circuit mechanisms accomplishing this feat are poorly understood. The general idea is that our brains use different active gating mechanisms for memory storage and memory systems whose dynamics evolve on vastly different timescales to consolidate memories for long-term storage. While some of these consolidation mechanisms are thought to take place at the system level, i.e., neuronal circuits or networks with different learning dynamics (McClelland et al., 1995; Schapiro et al., 2017), or dedicated circuits for replaying sensory experience (van de Ven et al., 2020), in this chapter, we focus on the putative synaptic mechanisms, typically referred to as “synaptic memory consolidation.”
@@ 3.3.1 The Ideal Observer Framework
The IOF simplifies the learning problem by focusing at individual synapses while ignoring all the complexity of the surrounding network and even the learning algorithm. Although a stark simplification, the IOF makes the learning problem theoretically more tractable and allows deriving fundamental insights about the effect of synaptic complexity on learning. Specifically, we assume that our memory system acquires new memories with the constant rate $r$. A constant rate of memory acquisition immediately links the memory capacity and the lifetime of a memory. Further, we assume that storing each memory is associated with synaptic changes that either depress or potentiate each synapse or leave it unchanged. In reality these plasticity events are generated by some learning algorithm, about which we only make few assumptions. We mainly require that the plasticity events are balanced, i.e., the up and down events cancel each other on average. This assumption is reasonable. Without it all synapses would soon saturate at their minimum or maximum value. Finally, we assume that each event is associated with storing a memory and that these memories are statistically independent.
@@ 3.3.2 Limits on memory lifetimes in simple synapses
a troubling result due to its logarithmic dependence (Fusi et al., 2005). Synapses are ideal loci of information storage in neuronal circuits because of their sheer abundance. However, the logarithmic dependence in Eq. (3) largely nullifies this advantage. The only choice we are left with for extending memory lifetimes is to set $q\ll 1$. Still, a small storage probability $q$ is also undesirable because it reduces our initial signal. Formally the above theoretical considerations underscore the fundamental character of the plasticity-stability dilemma. Further, the result shows that not even an idealized observer can leverage the sheer abundance of synapses ($N_{\mathrm{syn}}\gg 1$) to sidestep the problem. The above considerations raise the important question of how biological systems can store many memories with a large initial signal to noise ratio and safeguard them over extended periods of time.
@@ 3.3.3 Memory lifetimes in complex synapses
While the discrete state structure may be seen as a limitation of the model, Benna and Fusi (2016) later generalized the synaptic model to a system of interacting continuous variables with different interaction strengths (Fig. 5c), and showed that this could lead to similarly increased memory lifetime while also ensuring almost linear capacity scaling in the number of synapses, a significant improvement over previous models with $\sim\nicefrac{{1}}{{\sqrt{t}}}$ scaling, once more cementing the fundamental relevance of synaptic complexity for memory storage in neural networks.
The IOF outlined above demonstrates the potential of synaptic complexity and consolidation to mitigate the plasticity-stability dilemma. However, to enjoy power-law forgetting in an actual memory system requires working with real-world data, which violates some of the simplifying assumptions made above. It also requires a recall mechanism which may have to content itself with less than ideal observations of the system state. Modeling such a system is tricky because it requires dealing with high-dimensional data and a functional network model operating on these data, the latter of which is a major feat in itself. Here, deep learning has emerged as a practical tool for neuroscience that sidesteps the issue by providing a recipe for building complex neural networks using optimization principles (Richards et al., 2019).
@@ 4.1 Catastrophic forgetting
The struggle of ANNs to learn continually contrasts with our everyday experience of how we learn. We and animals tend to learn better when different tasks are learned independently. Moreover, we are capable of lifelong learning. Several circuit mechanisms have been proposed that may underlie this unique ability of biological brains (Parisi et al., 2019; Kudithipudi et al., 2022). First, the idea of complementary learning systems posits that different brain areas learn on separate timescales and make continual learning possible through their interactions (McClelland et al., 1995; Schapiro et al., 2017). The key idea is that one area, e.g., the hippocampus, acquires new memories quickly. While also susceptible to fast forgetting, it then copies this memory to a different brain area, e.g., the neocortex which learns and forgets more slowly. This avenue is also called system-level consolidation and may require some form of replay (van de Ven et al., 2020). It is covered elsewhere in this book. Second, synaptic consolidation requires synaptic complexity and is thought to accomplish a similar feat within a given neuronal circuit or brain area. This idea requires synapses and neurons to keep track of the information they store and then use this knowledge to inform synaptic plasticity to intelligently weave new memories into the neural memory fabric to minimize interference with previously stored information. We have already discussed phenomenological and theoretically motivated models that argued for synaptic complexity. Specifically, we looked at metaplasticity as prescribed by the Cascade model and how it can extend memory lifetimes. Thus far, the empirical success of applying the Cascade model to deep neural networks was limited, possibly because several of its underlying assumptions are violated by deep learning algorithms and real-world data. For instance, weight updates are not balanced, and the stored memories are not uncorrelated. Historically, synaptic theories of continual learning applied to ANNs thus took the idea of synaptic complexity but employed it differently as we will see next. We will cover the central ideas underlying recent continual learning algorithms that mitigate catastrophic forgetting in deep neural networks. After that, we will highlight two recent studies that have reinvigorated aspects of the original ideas of the Cascade model. But before discussing any of these topics, we must understand how ANNs learn from data by optimizing loss functions.
@@ 4.4 Synaptic importance measures
In elastic weight consolidation (EWC), Kirkpatrick et al. (2017) use the diagonal of the Fisher information metric as the importance measure $\Omega_{i}$. Specifically, this metric is the Hessian of the Kullback-Leibler divergence between the network parameters that minimize one task compared to the other. Or, in simpler terms, it can be understood as a measure of the local curvature of the loss landscape. That is, it assigns a high value to highly curved parameter space directions. Suppose we just finished training on Task 1. By definition, we are close to a minimum of $\mathcal{L}^{1}$ in parameter space. As we learn Task 2 next, we want to minimize parameter changes that change $\mathcal{L}^{1}$ because then we would make things worse on Task 1, i.e., we would forget. Thus, we want to make sure that we minimize our loss on Task 2 while ideally only walking into flat directions on $\mathcal{L}^{1}$. Flat means less curvature. To accomplish this, EWC more strongly penalizes updates in directions with high curvature on previous tasks. In practice, this strategy works remarkably well and has since become a standard for comparison.
@@ 4.6.1 Reinforcement learning with complex synapses
In another set of experiments, Kaplanis et al. (2018) investigated whether complex synapses could prevent within-task forgetting by alleviating the need for the replay dataset usually required to train deep RL agents. Considering the same tasks as above, the authors found for the Cart-Pole task that only the agent equipped with complex synapses could successfully learn online without the replay dataset, whereas vanilla agents were unable able to learn across a wide range of different learning rates. In contrast, the vanilla agents were able to learn the Catcher task without replay, which the authors explained by the more stationary experience distribution of the task. Overall, the authors concluded that synaptic metaplasticity can effectively reduce forgetting in ANNs without requiring an task-aware consolidation mechanism as in EWC. Still, it remains an open question how scalable this approach is and whether it transfers to more difficult deep RL problems like studied by (Kirkpatrick et al., 2017).
@@ 4.6.2 Synaptic metaplasticity in binarized neural networks
Laborieux et al. (2021) showed that the magnitude of the hidden weight naturally encodes an importance metric for the associated binarized weight, whereby higher hidden weight magnitude correspond to more important weights. The intuition is that if the loss is evaluated at the same binary weight, the hidden weights accumulate information about the gradient at this point without changing the sign (Fig. 13b). Based on this observation, Laborieux et al. (2021) designed a consolidation scheme inspired by the Cascade model (Fusi et al., 2005), in which depressing updates are multiplied by $f_{m}(w_{\rm h})$, which monotonically decreases with the hidden weight magnitude (Fig. 13c), thereby making it more difficult to decrease the absolute value of important synapses. The hyperparameter $m$ controls how fast weights are consolidated, and the case $m=0$ corresponds to the symmetric case. With such metaplasticity BNNs can continually learn sequences of tasks such as permuted MNIST, on par with EWC (Fig. 13d). Importantly, the consolidation mechanism is agnostic to transitions between tasks, and does not rely on an additional regularization loss. This feature prompted the authors to explore within-task forgetting in a streaming data setting in which different parts of the dataset were available at any give time (Fig. 13e). When learning Fashion MNIST or CIFAR-10 in this setting, the metaplastic network outperformed the control network without complex metaplastic synapses by a large margin, close to the performance level of joint training on complete dataset. Thus, metaplastic BNNs show that continual learning through synaptic consolidation can be achieved locally with complex synapses and agnostic to task boundaries in deep BNNs. It remains an exciting open question for future work to see whether and how similar principles can be extended to non-binary synapses.

### v2

@@ 1 Introduction
The vast neuronal networks that constitute our brain are the substrate underlying this remarkable ability of memory formation and storage. Memories are, in large part, stored as synaptic efficacy changes, i.e., the strength of the synaptic connections between neurons. Jointly, these changes cause lasting changes to the computational transformation that a given neuronal circuit performs, and it is this change that lies at the heart of memory. For instance, consider a neuronal circuit whose task is to associate a particular reward, e.g., some food, with a specific sensory stimulus. Upon experiencing a particular stimulus, e.g., a tone of a given frequency, the circuit in question may not produce any output initially. That is to say that it has not yet associated any value with this stimulus. Only once the tone is paired consistently with a reward does the brain circuit start producing an output specific to the paired stimulus. Thus, the next time the circuit receives the tone stimulus, its output will signal the value, e.g., the expected reward, even before the food has arrived. Meanwhile, the circuit response to another control stimulus, i.e., another tone, is unchanged — the circuit described above stores a memory of the rewarded stimulus. The paradigm described above is called Pavlovian conditioning, which honors Pavlov, who first performed such an experiment on his dog. And importantly, there are several brain areas, like, for instance, the amygdala, that are involved in such value learning.
@@
Many bio-chemical mechanisms exist that can change synaptic connections in an activity-dependent manner. These mechanisms are jointly called synaptic plasticity and are presumed to be the bio-physical basis of memory and learning. This book covers synaptic plasticity in-depth, so we won’t delve into them too deeply. For now, let us focus on the conceptual challenges these synaptic plasticity mechanisms must overcome to implement memory. There are several challenges. First, the synaptic changes must be stimulus-specific to form a distinct memory. Second, the changes need to be orchestrated so that the circuit does not succumb to pathological activity but remains in a healthy activity regime. Thus, plasticity also needs to ensure the stability of the circuit. This function is typically associated with various forms of homeostatic plasticity in neuronal circuits that act on different length- and timescales. Finally, plasticity must ensure that the circuit remains plastic for storing future memories without compromising the information already stored in the circuit.
@@
Our ability to learn continually is pervasive and intrinsically linked to how we learn. It applies to most learning problems requiring memory. Language is only one example. Other examples are physical skills such as riding a bicycle or playing volleyball. We must learn both skills in an often painstaking process, and they require overlapping sets of muscles, so at least some of the neuronal circuits involved are bound to be shared among them. We learn both tasks at different times, and yet learning to play volleyball does not impair our ability to ride a bike. A plethora of mechanisms enable us to learn continually. Some of them act at the system level, i.e., neuronal circuits or networks with different learning dynamics (McClelland et al., 1995; Schapiro et al., 2017; van de Ven et al., 2020), while others operate at the synaptic level. System level and replay approaches, also widely used in machine learning, are covered elsewhere in this book. This chapter focuses on the putative synaptic mechanisms.
In the following, we first review the classic Hopfield network, an influential associative memory model, which allows us to define essential concepts such as memory capacity and interference. The latter is linked to forgetting, and we will see how. Equipped with these concepts, we will explore several strands of research suggesting that synapses must possess an internal state or memory in their own right and intelligently use it to store new memories. To understand why, we will get to know the notion of memory lifetime and the fundamental requirement for memory consolidation mechanisms in the continual learning setting. Finally, we will turn to the more contemporary deep artificial neural network (ANN) models, the problem of catastrophic forgetting, and putative solutions based on synaptic consolidation and “intelligent” synaptic plasticity rules.
@@ 2 Learning and forgetting in Hopfield networks
which stores $p$ randomly and independently drawn binary patterns $\xi\in\{-1,1\}^{n}$ in which a one appears with a probability of 0.5. By convention neuron $i$ is active in memory pattern $\mu$ if $\xi_{i}^{\mu}=+1$. If we look at the above expression, we see that each term of the sum adds positive connections between neurons that are active (+1) in the same memory together. Conversely, if one neuron is active (+1) and the other inactive (-1) in a given pattern, the sign of the weight update is negative resulting in an inhibitory interaction between the two. Hence, we can think of learning in the Hopfield model as a Hebbian learning rule which defines assemblies of neurons with mutually excitatory connections as engrams for each stored memory (Hebb, 1949; cf. Fig. 1).
@@ 2.1 Memory capacity and forgetting
In the Hopfield model, the phenomenon is called “blackout catastrophe,” but it is also often referred to as “catastrophic forgetting” in the literature. Here, a word of caution is in order, as it is the first time we encounter the term catastrophic forgetting, and we will re-encounter it when discussing learning in ANNs (Section 4). We highlight it here because the term refers to fundamentally different phenomena in these two contexts. In the Hopfield model, catastrophic forgetting is linked to its finite memory capacity. It describes a phenomenon that impairs the network’s ability to recall any stored memory patterns through its dynamics. In the context of ANNs, as we will see later, it refers to old memories being quickly overwritten by new memories, a feature also called the palimpsest property. What adds to the confusion is that overwriting old memories through new memories is the solution to avoiding catastrophic forgetting in the Hopfield model. At the same time, forgetting is the primary problem for ANNs. While the whole meaning of this statement will only become clear at the end of this chapter, let us first understand why overwriting old memories solves the issue of catastrophic forgetting in the Hopfield network.
@@ 2.2 The stability-plasticity dilemma
It is called a paradox because the above statement contradicts our daily experience by which we constantly and quickly form new memories. For instance, we readily remember a person to whom we are introduced. But we also retain old memories, some of them for life. Our brains have evolved to deal with this plasticity-stability dilemma very efficiently. They learn continually by adding new memories with minimal impact on many old memories. The underlying synaptic and neuronal circuit mechanisms accomplishing this feat are poorly understood. The general idea is that our brains use different active gating mechanisms for memory storage and memory systems whose dynamics evolve on vastly different timescales to consolidate memories for long-term storage. While some of these consolidation mechanisms presumably act at the system level (McClelland et al., 1995; Schapiro et al., 2017), for instance, by actively replaying previously experienced activity patterns (Roxin and Fusi, 2013; van de Ven et al., 2020; Tomé et al., 2022), in this chapter, we focus on the putative synaptic mechanisms, typically referred to as “synaptic memory consolidation.”
@@ 3.3.1 The Ideal Observer Framework
The IOF simplifies the learning problem by focusing at individual synapses while ignoring all the complexity of the surrounding network and even the learning algorithm. Although a stark simplification, the IOF makes the learning problem theoretically more tractable and allows deriving fundamental insights about the effect of synaptic complexity on learning. Specifically, we assume that our memory system acquires new memories with the constant rate $r$. A constant rate of memory acquisition immediately links the memory capacity and the lifetime of a memory. Further, we assume that storing each memory is associated with synaptic changes that either depress or potentiate each synapse or leave it unchanged. In reality these plasticity events are generated by some learning algorithm, about which we only make few assumptions. We mainly require that the plasticity events are balanced, i.e., the up and down events cancel each other on average. This assumption is reasonable. Without it all synapses would soon saturate at their respective maximum or minimum, e.g., with an exponential distribution (Fusi and Abbott, 2007). Finally, we assume that each event is associated with storing a memory and that these memories are statistically independent.
@@ 3.3.2 Limits on memory lifetimes in simple synapses
a troubling result due to its logarithmic dependence (Amit and Fusi, 1994). Synapses are ideal loci of information storage in neuronal circuits because of their sheer abundance. However, the logarithmic dependence in Eq. (3) largely nullifies this advantage. The only choice we are left with for extending memory lifetimes is to set $q\ll 1$. Still, a small storage probability $q$ is also undesirable because it reduces our initial signal. Formally the above theoretical considerations underscore the fundamental character of the plasticity-stability dilemma. Further, the result shows that not even an idealized observer can leverage the sheer abundance of synapses ($N_{\mathrm{syn}}\gg 1$) to sidestep the problem. The above considerations raise the important question of how biological systems can store many memories with a large initial signal to noise ratio and safeguard them over extended periods of time.
@@ 3.3.3 Memory lifetimes in complex synapses
While the discrete state structure may be seen as a limitation of the model, Benna and Fusi (2016) later generalized the synaptic model to a system of interacting continuous variables with different interaction strengths (Fig. 5c), and showed that this could lead to similarly increased memory lifetime in which the SNR decreases as $\sim\nicefrac{{1}}{{\sqrt{t}}}$ while also ensuring almost linear capacity scaling in the number of synapses, a significant improvement over previous models with $\sim\nicefrac{{1}}{{t}}$ forgetting, once more cementing the fundamental relevance of synaptic complexity for memory storage in neural networks.
The IOF outlined above demonstrates the potential of synaptic complexity and consolidation to mitigate the plasticity-stability dilemma. However, to enjoy power-law forgetting in an actual memory system requires working with real-world data, which violates some of the simplifying assumptions made above. It also requires a recall mechanism which may have to content itself with less than ideal observations of the system state. While promising studies with a concrete recall mechanism exist that show the above ideas generalize to real-world data (Ji-An et al., 2023), modeling such systems remains challenging because it requires dealing with high-dimensional data and a functional network model operating on these data, the latter of which is a major feat in itself. Here, deep learning has emerged as a practical tool for neuroscience that sidesteps the issue by providing a recipe for building complex neural networks using optimization principles (Richards et al., 2019).
@@ 4.1 Catastrophic forgetting
The struggle of ANNs to learn continually contrasts with our everyday experience of how we learn. We and animals tend to learn better when different tasks are learned independently. Moreover, we are capable of lifelong learning. Several circuit mechanisms have been proposed that may underlie this unique ability of biological brains (Parisi et al., 2019; Kudithipudi et al., 2022). First, the idea of complementary learning systems posits that different brain areas learn on separate timescales and make continual learning possible through their interactions (McClelland et al., 1995; Schapiro et al., 2017). The key idea is that one area, e.g., the hippocampus, acquires new memories quickly. While also susceptible to fast forgetting, it then copies this memory to a different brain area, e.g., the neocortex which learns and forgets more slowly. This avenue is also called system-level consolidation and may require some form of replay (Roxin and Fusi, 2013; van de Ven et al., 2020; Tomé et al., 2022). It is covered elsewhere in this book. Second, synaptic consolidation requires synaptic complexity and is thought to accomplish a similar feat within a given neuronal circuit or brain area. This idea requires synapses and neurons to keep track of the information they store and then use this knowledge to inform synaptic plasticity to intelligently weave new memories into the neural memory fabric to minimize interference with previously stored information. We have already discussed phenomenological and theoretically motivated models that argued for synaptic complexity. Specifically, we looked at metaplasticity as prescribed by the Cascade model and how it can extend memory lifetimes. Thus far, the empirical success of applying the Cascade model to deep neural networks was limited, possibly because several of its underlying assumptions are violated by deep learning algorithms and real-world data. For instance, weight updates are not balanced, and the stored memories are not uncorrelated. Historically, synaptic theories of continual learning applied to ANNs thus took the idea of synaptic complexity but employed it differently as we will see next. We will cover the central ideas underlying recent continual learning algorithms that mitigate catastrophic forgetting in deep neural networks. After that, we will highlight two recent studies that have reinvigorated aspects of the original ideas of the Cascade model. But before discussing any of these topics, we must understand how ANNs learn from data by optimizing loss functions.
@@ 4.4 Synaptic importance measures
In elastic weight consolidation (EWC), Kirkpatrick et al. (2017) use the diagonal of the Fisher information metric as the importance measure $\Omega_{i}$. The Fisher information metric measures how much information a random variable carries about a parameter of interest. In the context of neural networks, it captures how sensitive the model’s predictions are to changes in the parameters. Close to a minimum the Fisher information metric is equivalent the Hessian of the loss. The Hessian is a square matrix of second-order partial derivatives that describes the curvature of the loss function. Thus EWC’s importance measure characterizes the local curvature of the loss landscape. That is, it assigns a high value to highly curved parameter space directions. Suppose we just finished training on Task 1. By definition, we are close to a minimum of $\mathcal{L}^{1}$ in parameter space. As we learn Task 2 next, we want to minimize parameter changes that change $\mathcal{L}^{1}$ because then we would make things worse on Task 1, i.e., we would forget. Thus, we want to make sure that we minimize our loss on Task 2 while ideally only walking into flat directions on $\mathcal{L}^{1}$. Flat means less curvature. To accomplish this, EWC more strongly penalizes updates in directions with high curvature on previous tasks. In practice, this strategy works remarkably well and has since become a standard for comparison.
@@ 4.6.1 Reinforcement learning with complex synapses
In another set of experiments, Kaplanis et al. (2018) investigated whether complex synapses could prevent within-task forgetting by alleviating the need for the replay dataset usually required to train deep RL agents. Considering the same tasks as above, the authors found for the Cart-Pole task that only the agent equipped with complex synapses could successfully learn online without the replay dataset, whereas vanilla agents were unable able to learn across a wide range of different learning rates. In contrast, the vanilla agents were able to learn the Catcher task without replay, which the authors explained by the more stationary experience distribution of the task. Overall, the authors concluded that synaptic metaplasticity can effectively reduce forgetting in ANNs without requiring an task-aware consolidation mechanism as in EWC. Still, it remains an open question how scalable this approach is and whether it transfers to more difficult deep RL problems like those studied by (Kirkpatrick et al., 2017).
@@ 4.6.2 Synaptic metaplasticity in binarized neural networks
Laborieux et al. (2021) showed that the magnitude of the hidden weight naturally encodes an importance metric for the associated binarized weight, whereby higher hidden weight magnitude correspond to more important weights. The intuition is that if the loss is evaluated at the same binary weight, the hidden weights accumulate information about the gradient at this point without changing the sign (Fig. 13b). Based on this observation, Laborieux et al. (2021) designed a consolidation scheme inspired by the Cascade model (Fusi et al., 2005), in which the depressing synaptic updates are multiplied by a metaplastic factor that decreases as a function of the hidden magnitude (Fig. 13c), thereby consolidating important synapses. With such metaplasticity BNNs can continually learn sequences of tasks such as permuted MNIST, on par with EWC (Fig. 13d). Importantly, the consolidation mechanism is agnostic to transitions between tasks, and does not rely on an additional regularization loss. This feature prompted the authors to explore within-task forgetting in a streaming data setting in which different parts of the dataset were available at any give time (Fig. 13e). When learning Fashion MNIST or CIFAR-10 in this setting, the metaplastic network outperformed the control network without complex metaplastic synapses by a large margin, close to the performance level of joint training on complete dataset. Thus, metaplastic BNNs show that continual learning through synaptic consolidation can be achieved locally with complex synapses and agnostic to task boundaries in deep BNNs. It remains an exciting open question for future work to see whether and how similar principles can be extended to non-binary synapses.